# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

### v1

@@ Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance. To address this, we propose a novel approach leveraging graph convolutional networks (GCNs), named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions. By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows GCNs to learn from the interconnected nature of language more effectively. Additionally, we explore the incorporation of phrase-based features to enhance classification accuracy, especially in low-resource settings. Our findings demonstrate that GCNs, augmented with these features, offer a promising solution for more accurate and context-aware question classification, bridging the gap between graph neural network research and practical educational applications.
**footnotetext: Equal contribution
Introduction
@@
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
Related Work
Automatic Question Classification
@@
One of the earliest works on Question Classification was in (Li and Roth 2002) which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features. Another reference work by (Huang, Thint, and Qin 2008) also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms. They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification (Hovy, Hermjakob, and Ravichandran 2002) via string matching in questions have also been explored in earlier works.
@@
Following these works, a review of automatic question classifiers from 2012 to 2017 (Silva, Bittencourt, and Maldonado 2019) revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches range from concatenating pre-trained word embeddings to form a question embedding (Godea and Nielsen 2018), to combining CNN and LSTM in a hybrid model (Zhou et al. 2016/12), effectively balancing feature extraction and sequence modeling, and was shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
Text Classification Using Text Graphs
@@
TextGCN (Yao, Mao, and Luo 2019) pioneered approaching text classification as a node classification problem. Yao et al. constructs a corpus-level graph with document nodes and word nodes—pointwise mutual information (PMI) as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output. Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation (Wu et al. 2019; Zhu and Koniusz 2021).
@@
Other works have also explored multi-graphs instead of a single corpus-level graph. TensorGCN (Liu et al. 2020) builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values. Each of these graphs has different word-word edge weights while sharing the same TF-IDF values for word-document edge weights, and they undergo intra- and inter-graph propagation. ME-GCN (Wang et al. 2022) constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings and similarity scores. SHINE (Wang et al. 2021) uses
@@
While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN (Wang, Han, and Poon 2022) extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
@@
Graph classification, in which you classify a document graph as opposed to a document node, has also been used as an inductive approach to text classification. Text-Level-GNN (Huang et al. 2019) uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair. Huang et al. also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighbourhood information via max-pooling and combining them via weighted sums, instead of convolution functions. Graph Fusion Network (Dai et al. 2022) constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs. GCN is applied to each subgraph, and the resulting output is concatenated, passed through an MLP layer, and average-pooled to obtain the final document embedding.
@@
There are also combinatory methods such as HyperGAT (Hypergraph Attention Networks) (Ding et al. 2020) building hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations.
Why Use Phrase Features?
@@
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks. We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. The effectiveness of phrases have been discussed in topic modeling for questions (Supraja, Khong, and Tatinati 2021; Supraja and Khong 2024), and this section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
@@
Unlike individual words, phrases (e.g., noun phrases and verb phrases) often encapsulate more specific and contextually relevant information. Consider the question ”How do enzymes function in the human body?” Here, the phrase ”enzymes function” (verb phrase) and ”human body” (noun phrase) are pivotal in identifying this as a biological process question, likely classified under ”Biology” or ”Physiology.” The verb phrase indicates the action or process in question, while the noun phrase specifies the domain of interest. In the question ”Describe the steps to solve a quadratic equation,” the phrases ”steps to solve” (verb phrase) and ”quadratic equation” (noun phrase) are linked by their relevance to problem-solving in mathematics.
This also applies to question analysis with part-of-speech (POS) tags. POS tags help in identifying the grammatical structure, but they do not capture the full semantic context. For example, knowing that ”role” is a noun does not reveal its specific relevance in a question about ”electric circuits.”
@@
By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question. For instance, in ”What are the causes of the French Revolution?”, the proximity and relationship between ”causes” and ”French Revolution” (via PPMI or other metrics) are key to categorizing this question under ”History” or ”Political Science.”
@@
Phrases provide additional context that can help disambiguate words that might have multiple meanings. For instance, in ”What is the role of the cell in an electric circuit?”, the noun phrase ”role of the cell” and its proximity to ”electric circuit” helps to clarify that ”cell” refers to a component of an electric circuit rather than a biological cell.
@@
Methodology
Feature Extraction and Graph Construction
@@
To construct meaningful graph representations from a corpus of question text, we selected several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table 1.
Word PPMI Word POS Phrase PPMI Phrase POS Named Entities
Node Type Word POS tag Phrase Phrase POS tag Named entity
Node Feature Word2Vec - PhraseBERT - TransE
Edge Weight PMI value PMI value PMI value PMI value Cosine similarity
Table 1: Summary of Extracted Features
Figure 1: Process flow of feature extraction, graph construction, and model training
@@
We first tokenize the given question corpus by whitespace to obtain a unique set of words. A word-level graph is constructed, with word nodes connected by edges weighted with PMI values, representing local co-occurrence relationships. Specifically, we utilize positive PMI (PPMI), given by max(PMI($word_{1},word_{2}$), 0). Pre-trained word embeddings from word2vec (Mikolov et al. 2013) are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
@@
POS tags for each word is obtained via the default set from NLTK ***https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
@@
Phrases are extracted from each question text via regex matching for verb phrases and noun phrases with spaCy ***https://spacy.io in Python, and a unique set of phrases in the corpus is obtained. Each question is represented as a sequence of verb phrases and noun phrases present in the text. Then a phrase-level graph (similar to the word-level graph) is constructed with PMI-weighted edges. Phrase embeddings are obtained from PhraseBERT (Wang, Thompson, and Iyyer 2021) to provide additional semantic information.
@@
During the phrase extraction process above, spaCy also provides word-level POS tags for each word in the extracted phrases. These word-level POS tags are concatenated to form phrase-level POS tags. For example, in the question ”Discuss the main objective of layout design rules”, using the regex (ADJ)(NOUN), the noun phrase [main objective] is extracted, together with the phrase-level POS tag of (ADJ_NOUN). Then, the PPMI values between each pair of phrase-level POS tags are calculated as edge weights for a phrase-level POS tag graph.
@@
Named entities (e.g., people, organizations, locations) carry significant meaning and often denote key elements within a question. A list of named entities is obtained from the NELL knowledge base (Carlson et al. 2010), and is used to extract named entities via string matching. With the short text nature of questions, it is difficult to obtain multiple entities from a single question to calculate co-occurrence statistics. Hence, we utilize TransE embeddings (Bordes et al. 2013) for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
@@
For all 5 graphs constructed above, the node features are initialized as one-hot vectors. By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
Model Metric NU ARC LREC Bloom TREC-coarse TREC-fine
CNN F1 0.085 0.226 0.198 0.752 0.782 0.627
Precision 0.048 0.171 0.141 0.763 0.783 0.653
Recall 0.333 0.333 0.333 0.750 0.790 0.644
Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653 0.102
Precision 0.606 0.554 0.511 0.465 0.692 0.086
Recall 0.620 0.598 0.495 0.422 0.642 0.109
TextGCN (Yao, Mao, and Luo 2019) F1 0.722 0.694 0.671 0.663 0.730 0.505
Precision 0.715 0.687 0.677 0.652 0.675 0.500
Recall 0.735 0.735 0.679 0.680 0.780 0.589
Text-Level-GNN (Huang et al. 2019) F1 0.185 0.200 0.404 0.287 0.623 0.186
Precision 0.694 0.362 0.424 0.250 0.651 0.216
Recall 0.183 0.275 0.424 0.344 0.606 0.184
HyperGAT (Ding et al. 2020) F1 0.715 0.372 0.745 0.173 0.678 0.579
Precision 0.733 0.386 0.766 0.176 0.673 0.590
Recall 0.703 0.374 0.737 0.178 0.691 0.626
TensorGCN (Liu et al. 2020) F1 0.748 0.700 0.717 0.114 0.140 0.005
Precision 0.742 0.690 0.716 0.184 0.176 0.025
Recall 0.757 0.716 0.719 0.185 0.145 0.071
SHINE (Wang et al. 2021) F1 0.560 0.610 0.620 0.459 0.560 0.197
Precision 0.553 0.600 0.620 0.461 0.568 0.208
Recall 0.583 0.623 0.627 0.473 0.613 0.242
ME-GCN (Wang et al. 2022) F1 0.632 0.603 0.601 0.607 0.659 0.669
Precision 0.738 0.601 0.607 0.629 0.718 0.729
Recall 0.619 0.620 0.608 0.592 0.634 0.642
InducT-GCN (Wang, Han, and Poon 2022) F1 0.661 0.667 0.649 0.533 0.688 0.165
Precision 0.738 0.649 0.648 0.754 0.753 0.209
Recall 0.639 0.711 0.654 0.478 0.671 0.173
PQ-GCN F1 0.724 0.712 0.751 0.672 0.801 0.201
Precision 0.723 0.695 0.754 0.692 0.882 0.224
Recall 0.727 0.750 0.749 0.662 0.777 0.257
@@
Table 2: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
Model Architecture

### v2

@@ Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance. To address this, we propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions. By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows the model to learn from the interconnected nature of language more effectively. Additionally, we explore the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics. Our findings demonstrate that the proposed model, augmented with these features, offer a promising solution for more robust and context-aware question classification, bridging the gap between graph neural network research and practical educational applications of AI.
Index Terms:
natural language processing, graph convolutional network, question classification
11footnotetext: Equal contribution
I Introduction
@@
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [1] and [2] which are reviewed in Section II-A. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
II Related Work
II-A Automatic Question Classification
@@
One of the earliest works on Question Classification was in [3] which made use of support vector machines (SVMs) and explored a variety of features like lexical, synctactic, and semantic features. Another reference work by [4] also discusses usage of head words and hypernyms as features for SVM and maximum entropy (ME) algorithms. They demonstrated the effectiveness of statistical learning methods in classifying questions, while rule-based classification [5] via string matching in questions have also been explored in earlier works.
@@
Following these works, a review of automatic question classifiers from 2012 to 2017 [6] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches include concatenating pre-trained word embeddings to form a question embedding [7], incorporating subword information in character-based convolutional neural network (CNN) [8], and combining CNN and long short-term memory (LSTM) in a hybrid model [9]. These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN. The first work [1] incorporates dependency parsing between words, pointwise mutual information (PMI) between regular expressions, cosine similarity between phrase embeddings, and KL divergence between topics in a heterogeneous graph for question and short document classification. The authors based the architecture on TensorGCN [10] and 4 heterogeneous node sets and adjacency matrices, while PQ-GCN is based on SHINE [11] and uses 5 homogeneous node sets and adjacency matrices which reduces sparsity in input data. The second work [2] uses word graph based on dependency parsing, and trains a graph attention network (GAT) [12]. This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of a question, given their short lengths.
II-B Text Classification Using Text Graphs
@@
TextGCN [13] pioneered approaching text classification as a node classification problem. Yao et al. constructs a corpus-level graph with document nodes and word nodes—PMI as word-word edge weights and text frequency-inverse document frequency (TF-IDF) as word-document edge weights—and applies a 2-layer GCN to obtain the class output. Several works propose similar graph construction and node initialization based on TextGCN, with more efficient or better regularized graph propagation [14, 15].
@@
Other works have also explored multi-graphs instead of a single corpus-level graph. TensorGCN [10] builds a tensor graph from three separate graphs: semantic graph via word embeddings, syntactic graph via grammar parsing, and sequential graph via PMI values. Each of these graphs has different word-word edge weights while sharing the same TF-IDF values for word-document edge weights, and they undergo intra- and inter-graph propagation. ME-GCN [16] constructs a graph that has multi-dimensional word-word, word-document, and document-document edges, using Word2Vec and Doc2Vec embeddings for node features, and normalized embedding distance and TF-IDF for edge features. SHINE [11] uses word graphs based on three word-level relationships—PMI, POS tag co-occurrence, and cosine similarities between embeddings of entity pairs appearing in the corpus—which are easy to obtain at a low cost using common natural language processing techniques.
@@
While node classification is a transductive approach, where one needs to build a corpus-level graph consisting both train and test document nodes, InducT-GCN [17] extends the task into a inductive approach, where a training corpus graph is constructed first similarly to TextGCN, and any new test corpus is built as a virtual subgraph for inference.
@@
Graph classification, in which you classify a document graph as opposed to a document node, has also been used as an inductive approach to text classification. Text-Level-GNN [18] uses a sliding window to build multiple graphs, each with a small number of nodes and edges, with trainable edge weights that are shared across the graphs for the same word pair. Huang et al. also uses a message passing mechanism instead of a GCN architecture, relying on aggregating neighborhood information via max-pooling and combining them via weighted sums, instead of convolution functions. Graph Fusion Network [19] constructs four individual global word co-occurrence graphs, and each document is represented as with a set of four subgraphs. GCN is applied to each subgraph, and the resulting output is concatenated, passed through an MLP layer, and average-pooled to obtain the final document embedding.
@@
There are also combinatory methods based on GAT. HyperGAT (Hypergraph Attention Networks) [20] builds hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations. Contrastive learning has also been utilized in conjunction with topic modeling and GAT to improve classification performance on short texts [21].
III Why Use Phrase Features?
@@
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks. We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. The effectiveness of phrases in providing deeper semantic connections in the text has been discussed in topic modeling for questions [22, 1], as well as in graph-based document filtering tasks [23]. This section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
Figure 1: Corpus-level graph construction with extracted features
@@
Unlike individual words, phrases (e.g. noun phrases and verb phrases) often encapsulate more specific and contextually relevant information. Consider the question “How do enzymes function in the human body?” The phrase “enzymes function” and “human body” are pivotal in identifying this as a biological process question, likely classified under “Biology” or “Physiology.” The phrases indicates the action or process in question, and also specifies the domain of interest. In the question “Describe the steps to solve a quadratic equation,” the phrases “steps to solve” and “quadratic equation” are linked by their relevance to problem-solving in mathematics. This also applies to question analysis with part-of-speech (POS) tags. POS tags help in identifying the grammatical structure, but they do not capture the full semantic context when observed individually.
@@
By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question. For instance, in “What are the primary causes of the French Revolution?”, the proximity and relationship between “primary causes” and “French Revolution” are key to categorizing this question under “History” or “Political Science.”
@@
Phrases provide additional context that can help disambiguate words that might have multiple meanings. For instance, in “What is the role of the cell in an electric circuit?”, the phrase “role of the cell” and its proximity to “electric circuit” helps to clarify that “cell” refers to a component of an electric circuit rather than a biological cell.
@@
IV Methodology
IV-A Feature Extraction and Graph Construction
@@
To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table I.
TABLE I: Summary of Extracted Features
Node Type Node Feature Edge Weight
Word PPMI Word Word2Vec PMI value
Word POS POS tag - PMI value
Phrase PPMI Phrase PhraseBERT PMI value
Phrase POS Phrase POS tag - PMI value
Named Entities Named entity TransE Cosine similarity
Figure 2: Overview of model architecture
@@
We first tokenize the given question corpus by whitespace to obtain a unique set of words. A word-level graph is constructed, with word nodes connected by edges weighted with PMI values, representing local co-occurrence relationships. Specifically, we utilize positive PMI (PPMI) with a window size of 5, given by max(PMI($word_{1},word_{2}$), 0). PMI is given by:
$\text{PMI}(w_{1},w_{2})=\log\frac{P(w_{1},w_{2})}{P(w_{1})P(w_{2})}$ (1)
where $P(w_{1})=\frac{\#W(w_{1})}{\#W}$, $P(w_{2})=\frac{\#W(w_{2})}{\#W}$, $P(w_{1},w_{2})=\frac{\#W(w_{1},w_{2})}{\#W}$. #$\#W(w)$ is the number of sliding windows that a word $w$ in the corpus, and #$\#W(w_{1},w_{2})$ is the number of sliding windows that words $w_{1}$ and $w_{2}$ appear together in the corpus. #$\#W$ is the total number of sliding windows in the corpus.
Pre-trained word embeddings from word2vec [24] are also obtained for each word in the corpus, to be used as additional semantic information by concatenating with node embeddings.
TABLE II: PQ-GCN Model Parameters
Word Embedding Size 300
Phrase Embedding Size 768
Entity Embedding Size 100
Sliding Window Size for PMI 5
Hidden Dimension Size 200
No. of Layers in GCNs 2
Dropout 0.7
Threshold for Question Cosine Similarity 2.7
Optimizer Adam
Learning Rate 1e-3
Weight Decay 1e-4
@@
POS tags for each word is obtained via the default set from NLTK  [^1] https://nltk.org and PPMI values between each pair of POS tags are calculated as edge weights for a POS tag graph.
@@
Phrases are extracted from each question text via POS tag regex matching for noun and verb phrases with spaCy  [^2] https://spacy.io in Python, and a set of unique phrases in the corpus is obtained. The regex patterns for noun and verb phrases are given by (2) and (3) respectively:
[…]
$(\langle\text{NOUN}\rangle|\langle\text{PROPN}\rangle\langle\text{PART}\rangle?)+$ (2)
$\mu_{V}=$ $\langle\text{AUX}\rangle*\langle\text{ADV}\rangle*\langle\text{VERB}\rangle$ (3)
where ????, *, and $+$ represent zero or one, zero or more, and one or more occurrences of the preceding POS tag respectively.
Each question is represented as a sequence of verb phrases and noun phrases present in the text. Then a phrase-level graph (similar to the word-level graph) is constructed with PMI-weighted edges. Phrase embeddings are obtained from PhraseBERT [25] to provide additional semantic information.
TABLE III: Classification Label Distributions
Dataset NU ARC LREC Bloom TREC
Dataset Size 596 279 344 2522 5952
Classification Cognitive complexities Reasoning capabilities Expected answer types Educational objectives Question topics
Avg. # of Words 10.8 24.8 16.1 14.8 10.0
Class Labels Transfer (49.7%) Linguistic matching Very short answer (41.6%) Comprehension (38.1%) Entity (22.6%)
Knowledge (33.5%) (47.0%) Context sensitive (36.0%) Knowledge (13.6%) Description and abstract
Application (16.8%) Basic facts (28.0%) Answers will vary (22.4%) Evaluation (12.6%) concept (21.8%)
Hypothetical (25.0%) Application (12.5%) Human being (21.6%)
Analysis (12.1%) Numeric value (17.0%)
Synthesis (11.1%) Location (15.4%)
Abbreviation (1.6%)
TABLE IV: CNN Model Parameters
Parameter Layer 1 Layer 2 Layer 3 Layer 4
Type Conv2d Conv2d Conv2d Linear
Embedding Size 300
Sequence Length 15
Number of Filters 100 100 100
Kernel Size 3x3 4x4 5x5
Stride 1x1 1x1 1x1
Activation Function Softmax
Dropout Rate 0.5
TABLE V: Bi-LSTM Model Parameters
Parameter Layer 1 Layer 2 Layer 3
Type Bi-LSTM Bi-LSTM Linear
Embedding Size 300
Hidden Dimension Size 128 128
Activation Function Softmax
@@
[…]
@@
Named entities (e.g., people, organizations, locations) carry significant meaning and often denote key elements within a question. A list of named entities is obtained from the NELL knowledge base [26], and is used to extract named entities via string matching. With the short text nature of questions, it is difficult to obtain multiple entities from a single question to calculate co-occurrence statistics. Hence, we utilize TransE embeddings [27] for each named entity and calculate the cosine similarity between embeddings of each pair of named entities, to be used as edge weights in constructing a named entity graph.
@@
From the corpus, the above features are extracted and the respective relationships are computed to obtain edge weights, which are then used to construct the graph for each feature. The examples of graphs constructed are shown in Figure 1. For the corpus-level graphs constructed above, the node features are initialized as one-hot vectors except for entity graph, which uses the TransE embeddings directly as node features. The mapping of which text features are present in which question is used to generate question-level embeddings during propagation. By combining these diverse feature extraction methods, we construct a comprehensive and multi-faceted graph representation of the question text, capturing various semantic, syntactic, and named entity information crucial for question classification.
@@
TABLE VI: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
Best F1, precision, and recall scores are in bold and the second best scores are underlined for each dataset.
Model Metric NU ARC LREC Bloom TREC
CNN F1 0.085 0.226 0.198 0.752 0.782
Precision 0.048 0.171 0.141 0.763 0.783
Recall 0.333 0.333 0.333 0.750 0.790
Bi-LSTM F1 0.607 0.564 0.480 0.425 0.653
Precision 0.606 0.554 0.511 0.465 0.692
Recall 0.620 0.598 0.495 0.422 0.642
TextGCN [13] F1 0.722 0.694 0.671 0.663 0.730
Precision 0.715 0.687 0.677 0.652 0.675
Recall 0.735 0.735 0.679 0.680 0.780
Text-Level-GNN [18] F1 0.185 0.200 0.404 0.287 0.623
Precision 0.694 0.362 0.424 0.250 0.651
Recall 0.183 0.275 0.424 0.344 0.606
HyperGAT [20] F1 0.715 0.372 0.745 0.173 0.678
Precision 0.733 0.386 0.766 0.176 0.673
Recall 0.703 0.374 0.737 0.178 0.691
TensorGCN [10] F1 0.412 0.499 0.566 0.107 0.805
Precision 0.421 0.557 0.692 0.147 0.851
Recall 0.450 0.604 0.575 0.157 0.775
SHINE [11] F1 0.560 0.610 0.620 0.459 0.560
Precision 0.553 0.600 0.620 0.461 0.568
Recall 0.583 0.623 0.627 0.473 0.613
ME-GCN [16] F1 0.632 0.603 0.601 0.607 0.659
Precision 0.738 0.601 0.607 0.629 0.718
Recall 0.619 0.620 0.608 0.592 0.634
InducT-GCN [17] F1 0.661 0.667 0.649 0.533 0.688
Precision 0.738 0.649 0.648 0.754 0.753
Recall 0.639 0.711 0.654 0.478 0.671
PQ-GCN F1 0.724 0.712 0.751 0.672 0.801
Precision 0.723 0.695 0.754 0.692 0.882
Recall 0.727 0.750 0.749 0.662 0.777
IV-B Model Architecture